# Birational automorphism groups of Severi–Brauer surfaces over the field of rational numbers

### v2

@@ Abstract.
We prove that the only non-trivial finite subgroups of birational automorphism group of non-trivial Severi–Brauer surfaces over the field of rational numbers are $\mathbb{Z}/3\mathbb{Z}$ and $(\mathbb{Z}/3\mathbb{Z})^{2}.$ Moreover, we show that $(\mathbb{Z}/3\mathbb{Z})^{2}$ is contained in $\mathrm{Bir}(V)$ for any Severi–Brauer surface $V$ over a field of characteristic zero, and $(\mathbb{Z}/3\mathbb{Z})^{3}$ is contained in $\mathrm{Bir}(V)$ for any Severi–Brauer surface $V$ over a field of characteristic zero which contains a non-trivial cube root of unity. All the results of the paper hold for perfect fields of characteristic different from $2$ and $3$
@@ 1. Introduction
Like $\mathrm{Cr}_{n}(\mathbf{F}),$ the group $\mathrm{Bir}(V)$ of birational automorphisms of a Severi–Brauer variety $V$ also has a complicated structure (cf. [11] and [22]). A classification of finite groups that appear as subgroups of $\mathrm{Bir}(V)$ for non-trivial Severi–Brauer surfaces over various fields of characteristic zero was given in [18, Theorem 1.3]. On the other hand, there is no simple way to decide which of them are realized for a given field, or for a given Severi–Brauer surface. Meanwhile, we have a very simple description of finite subgroups of $\mathrm{Bir}(V)$ over a field of characteristic different from $3$ which contains no roots of unity of any odd degree.
Theorem 1.2 ([18, Corollary 1.4 and Remark 7.2]).
Let $V$ be a non-trivial Severi–Brauer surface over a field of characteristic different from $3$ which contains no roots of unity of any odd degree and let $G\subset\mathrm{Bir}(V)$ be a finite subgroup. Then we have $G\subset(\mathbb{Z}/3\mathbb{Z})^{3}.$
@@
Let $V$ be a non-trivial Severi–Brauer surface over a field of characteristic different from $2$ and $3$ which contains no roots of unity of any odd degree, and let $G$ be a finite group. Then $G$ is isomorphic to a subgroup of $\mathrm{Bir}(V)$ if and only if $G\subset(\mathbb{Z}/3\mathbb{Z})^{2},$ and $G$ is isomorphic to a subgroup of $\mathrm{Aut}(V)$ if and only if $G\subset\mathbb{Z}/3\mathbb{Z}.$
As a corollary we have the following result about finite subgroups of birational automorphism group of a Severi–Brauer surface over the field $\mathbb{Q}$ of rational numbers.
Corollary 1.4.
@@
Let $V$ be a Severi–Brauer surface over a field $\mathbf{F}$ of characteristic different from $2$ and $3$ Then
@@
This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265). The author was partially supported by the HSE University Basic Research Program, the Moebius Contest Foundation for Young Scientists and by by the BASIS foundation grant ”Young Russian Mathematics”.
@@ 4. Cubic surfaces
It remains to check that $r$ lies in $\mathrm{Aut}(S).$ The element $r$ commutes with the element $b,$ and therefore, by Lemma 4.9 it lies in the centralizer of $b$ in $\mathrm{Aut}(S_{\overline{\mathbf{F}}})\subset W(\mathrm{E}_{6}).$ Moreover, since the element $r$ commutes with the image of the Galois group in $W(\mathrm{E}_{6}),$ by Lemma 4.3 it is contained in $\mathrm{Aut}(S).$
@@
In this section we prove Proposition 1.6. First of all, let us study the $3$-groups in the automorphism groups of del Pezzo surfaces and conic bundles. Recall that the degree of a del Pezzo surface $S$ is defined as $d=(K_{S})^{2}.$ The following fact is well-known.
@@
Let $\mathbf{F}$ be a field of characteristic different from $3$ which does not contain non-trivial cube roots of unity. Let $V$ be a Severi–Brauer surface over $\mathbf{F}$. Then
@@
Assume that $G\simeq(\mathbb{Z}/3\mathbb{Z})^{3}$ is contained in $\mathrm{Bir}(V).$ Then $G$ acts biregularly either on a del Pezzo surface, or on a conic bundle $\phi:S\to B$ over a geometrically rational curve $B$ such that $\phi$ is equivariant with respect to $G$ (see, for example, [5, Lemma 3.6] and [15, Lemma 14.1.1]). By Lemma 5.3 the latter case is impossible. By Lemma 5.2 the group $G$ does not act biregularly on a del Pezzo surface of degree $d\neq 3.$ Finally, by Lemma 4.1 the group $G$ does not act biregularly on a del Pezzo surface of degree $d=3.$

### v4

@@ Abstract.
We prove that the only non-trivial finite subgroups of birational automorphism group of non-trivial Severi–Brauer surfaces over the field of rational numbers are $\mathbb{Z}/3\mathbb{Z}$ and $(\mathbb{Z}/3\mathbb{Z})^{2}.$ Moreover, we show that $(\mathbb{Z}/3\mathbb{Z})^{2}$ is contained in $\mathrm{Bir}(V)$ for any Severi–Brauer surface $V$ over a field of characteristic different from $2$ and $3$, and $(\mathbb{Z}/3\mathbb{Z})^{3}$ is contained in $\mathrm{Bir}(V)$ for any Severi–Brauer surface $V$ over a field of characteristic different from $2$ and $3$ which contains a non-trivial cube root of unity.
@@ 1. Introduction
Like $\mathrm{Cr}_{n}(\mathbf{F}),$ the group $\mathrm{Bir}(V)$ of birational automorphisms of a Severi–Brauer variety $V$ also has a complicated structure (cf. [11] and [22]). A classification of finite groups that appear as subgroups of $\mathrm{Bir}(V)$ for non-trivial Severi–Brauer surfaces over various fields of characteristic zero was given in [18, Theorem 1.3]. On the other hand, there is no simple way to decide which of them are realized for a given field, or for a given Severi–Brauer surface.
Theorem 1.2 ([18, Corollary 1.4]).
Let $V$ be a non-trivial Severi–Brauer surface over $\mathbb{Q}$ and let $G\subset\mathrm{Bir}(V)$ be a finite subgroup. Then we have $G\subset(\mathbb{Z}/3\mathbb{Z})^{3}.$
@@
Let $V$ be a Severi–Brauer surface over a perfect field $\mathbf{F}$ of characteristic different from $2$ and $3$ Then
@@
This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265). The author was partially supported the Moebius Contest Foundation for Young Scientists and by the BASIS foundation grant ”Young Russian Mathematics”.
@@ 4. Cubic surfaces
It remains to check that $r$ lies in $\mathrm{Aut}(S).$ The element $r$ commutes with the element $b,$ and therefore, by Lemma 4.9 it lies in the centralizer of $b$ in the automorphism group $\mathrm{Aut}(S_{\overline{\mathbf{F}}})\subset W(\mathrm{E}_{6}).$ Moreover, since the element $r$ commutes with the image of the Galois group in $W(\mathrm{E}_{6}),$ by Lemma 4.3 it is contained in $\mathrm{Aut}(S).$
@@
In this section we prove Proposition 1.5. First of all, let us study the $3$-groups in the automorphism groups of del Pezzo surfaces and conic bundles. Recall that a smooth projective surface $S$ is a del Pezzo surface if its anticanonical class $-K_{S}$ is ample. The degree of a del Pezzo surface $S$ is defined as $d=(K_{S})^{2}.$ The following fact is well-known.
@@
Let $\mathbf{F}$ be a perfect field of characteristic different from $3$ which does not contain non-trivial cube roots of unity. Let $V$ be a Severi–Brauer surface over $\mathbf{F}$. Then
@@
Assume that $G\simeq(\mathbb{Z}/3\mathbb{Z})^{3}$ is contained in $\mathrm{Bir}(V).$ Since $\mathbf{F}$ is perfect, the group $G$ acts biregularly either on a del Pezzo surface, or on a conic bundle $\phi:S\to B$ over a geometrically rational curve $B$ such that $\phi$ is equivariant with respect to $G$ (see, for example, [5, Lemma 3.6] and [15, Lemma 14.1.1]). By Lemma 5.3 the latter case is impossible. By Lemma 5.2 the group $G$ does not act biregularly on a del Pezzo surface of degree $d\neq 3.$ Finally, by Lemma 4.1 the group $G$ does not act biregularly on a del Pezzo surface of degree $d=3.$